# SafeVLA: Towards Safety Alignment of Vision- Language-Action Model via Constrained Learning

### v2

@@ Abstract
Vision-language-action models (VLAs) show potential as generalist robot policies. However, these models pose extreme safety challenges during real-world deployment, including the risk of harm to the environment, the robot itself, and humans. How can safety constraints be explicitly integrated into VLAs? We address this by exploring an integrated safety approach (ISA), systematically modeling safety requirements, then actively eliciting diverse unsafe behaviors, effectively constraining VLA policies via safe reinforcement learning, and rigorously assuring their safety through targeted evaluations. Leveraging the constrained Markov decision process (CMDP) paradigm, ISA optimizes VLAs from a min-max perspective against elicited safety risks. Thus, policies aligned through this comprehensive approach achieve the following key features: (I) effective safety-performance trade-offs, this exploration yields an 83.58% safety improvement compared to the current state-of-the-art method, while also maintaining task performance (+3.85%). (II) strong safety assurance, with the ability to mitigate long-tail risks and handle extreme failure scenarios. (III) robust generalization of learned safety behaviors to various out-of-distribution perturbations. Our data, models and newly proposed benchmark environment are available at https://pku-safevla.github.io.
@@ 1 Introduction
Embodied AI aims to develop a generalist policy that can perform perception, interaction, reasoning, and adaptation in the physical world liu2024aligning . Building on the emergence of large language models (LLMs) and vision-language models (VLMs), vision-language-action models (VLAs) brohan2022rt ; o2023open ; team2024octo ; kim2024openvla advance this field by enabling robots to follow vision-language instructions and perform tasks in real-world environments. As these models continue to evolve, they have the potential to become generalist robot policies reed2022generalist ; ma2024survey , capable of executing previously unseen instructions and effectively generalizing behaviors across a diverse range of robot embodiments, scenes, skills, and objects o2023open . Ensuring the alignment of these models with human values and safety has become more critical than ever kaddour2023challenges ; ji2023ai , due to their increasing complexity and power dubey2024llama ; openai2024o1 ; liu2024deepseek . While significant progress has been made in task performance, the explicit integration of safety mechanisms remains an open challenge.
@@
The safety risks of LLMs and VLMs have been extensively studied, with existing methods such as data augmentation gulcehre2023reinforced , content moderation inan2023llama ; chi2024llama , reinforcement learning from human feedback (RLHF) ouyang2022training ; touvron2023llama , Safe-RLHF dai2023safe , and lightweight alignment ji2024aligner . However, these safety mechanisms cannot be directly applied to VLAs, as there is a substantial gap between the abstract safety concerns at the model intention level and the unique safety challenges posed by the complex and unpredictable physical world guiochet2017safety . Despite large-scale behavior cloning and careful alignment in existing VLAs hu2024flare ; zhang2024grape , the most advanced models have yet to explicitly define and integrate safety as an integral aspect of their design brohan2023rt ; gu2023rt ; ehsani2024spoc ; belkhale2024rt ; black2024pi0 ; liu2024rdt . This fundamental limitation motivates an urgent need to explore methodologies capable of explicitly embedding safety constraints into the VLAs zacharaki2020safety ; falco2021governing .
@@ 2 Related Work
Vision-Language-Action Models.  Vision-language-action models (VLAs) brohan2022rt ; brohan2023rt ; o2023open ; ehsani2024spoc ; team2024octo ; kim2024openvla ; black2024pi0 represent a significant step towards generalist robots capable of executing complex tasks based on multimodal instructions in diverse environments reed2022generalist ; ma2024survey . These models, often built upon powerful foundation models zawalski2024robotic ; liu2024rdt ; pertsch2025fast and trained on large-scale trajectory datasets o2023open , demonstrate impressive task performance and generalization ability wang2024towards . However, their real-world deployment is hindered by safety concerns inherent to physical interaction guiochet2017safety ; zacharaki2020safety . While safety alignment is actively researched for LLMs and VLMs ouyang2022training ; bai2022constitutional ; dai2023safe ; ji2024beavertails ; ji2024aligner , methods focusing on mitigating abstract risks like harmful content generation ganguli2022red ; hendrycks2023overview do not readily address the concrete physical hazards faced by embodied agents. Current VLA training, typically relying on imitation learning (IL) ehsani2024spoc or standard reinforcement learning (RL) fine-tuning hu2024flare ; zhang2024grape , lacks mechanisms for explicitly integrating and enforcing safety constraints, leaving a critical gap for reliable deployment falco2021governing .
Safe Reinforcement Learning.  Safe reinforcement learning (SafeRL) within the constrained Markov decision process (CMDP) framework altman2021constrained ; gu2022review , offers a formal approach to policy optimization where an agent learns to maximize task rewards while explicitly satisfying predefined safety constraints. This paradigm contrasts with heuristic methods like reward shaping, which indirectly encode safety preferences and lack formal guarantees. While SafeRL techniques have been explored for aligning foundation models (e.g., Safe-RLHF dai2023safe ), applying them to high-dimensional, multimodal VLAs operating in complex physical environments poses unique challenges act2024artificial . Model-free, first-order optimization methods compatible with the CMDP formulation, such as Lagrangian-based approaches PIDLag2020 ; dai2023augmented , are promising for VLAs as they avoid restrictive assumptions about system dynamics or state structure, making them suitable for learning from raw perceptual inputs like RGB images ji2024omnisafe . Our work systematically explores the application of these principles to VLA safety alignment.
@@ 3 Problem Formulation
Constrained Markov Decision Process.  The constrained Markov decision process (CMDP) (altman2021constrained, ) is commonly used to model dynamic decision-making under uncertainty when multiple objectives are present. In this framework, the policy aims to maximize one objective while satisfying constraints on the others. A CMDP is defined as a tuple $(\mathcal{S},\mathcal{A},\mathbb{P},r,\mathcal{C},\mu,\gamma)$, where $\mathcal{S}$ is state space, $\mathcal{A}$ is action space. $\mathbb{P}(s^{{}^{\prime}}|s,a)$ is probability of state transition from $s$ to $s^{{}^{\prime}}$ after playing $a$. $r(\cdot):\mathcal{S}\times\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$, and $r(s^{{}^{\prime}}|s,a)$ denotes the reward that the agent observes when state transition from $s$ to $s^{{}^{\prime}}$ after it plays $a$. The set $\mathcal{C}=\{(c_{i},b_{i})\}_{i=1}^{m}$, where $c_{i}$ are cost functions: $c_{i}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$, and limits are $b_{i}$, $i=1,\cdot,m$. $\mu(\cdot):\mathcal{S}\rightarrow[0,1]$ is the initial state distribution and $\gamma\in(0,1)$. Let $\mathcal{H}_{t}$ be the set of all possible trajectories $(s_{0},a_{0},\dots,s_{t-2},a_{t-2},s_{t-1})$ of length $t$.
@@ 4 Implementing the Integrated Safety Approach
We argue that VLA safety requires an integrated safety approach (ISA), rather than a single method. Specifically, an ISA address four interconnected aspects, as shown in Figure 1: (i) modeling safety-critical aspects of tasks and environments; (ii) eliciting latent and diverse unsafe behaviors from existing policies; (iii) constraining the VLA’s learning process to integrate these safety considerations; and (iv) assuring the resulting model’s safety through rigorous and targeted evaluation. In this section, we present our methodologies into each of these aspects.
@@ 4.3 Constraining Policies: Safe Reinforcement Learning for Alignment
A preliminary step is translating safety predicates ($\phi,\psi$) into cost signals for the cost-returns $\mathcal{J}_{c_{i}}(\bm{\theta})$. State-action predicate ($\phi_{k}$) violations incur a cost of 1 at the violating timestep $t$, otherwise 0. For trajectory-level predicates ($\psi_{j}$), a cost of 1 is attributed solely to the final step of the violating segment in this initial exploration. The credit assignment for $\psi_{j}$ remains an area for exploration in future work. The Lagrangian method is a general solution for SafeRL. By employing the Lagrangian relaxation technique (NoceWrig06, ), Equation 2 is transformed into an unconstrained safe optimization problem:
@@ 4.4 Safety Assurance: Evaluating Aligned VLAs
Type Methods SR $\uparrow$ CC $\downarrow$ SR $\uparrow$ CC $\downarrow$ SR $\uparrow$ CC $\downarrow$
@@ 5.1 Experimental Setup
Tasks and Environments.  Our primary experiments utilize Safety-CHORES. To contextualize the unique challenges posed by Safety-CHORES, we also conduct comparisons on other benchmarks kolve2017ai2 ; deitke2020robothor ; deitke2022️ focusing on object navigation and generally lack the safety features of Safety-CHORES.
@@ 5.2.2 Qualitative Insights: Risk Handling and Failure Modes
Figure 5: Comparative performance of VLA models on multiple benchmarks. Left: SR of each model per benchmark. Right: CC incurred by each model on these benchmarks.
@@ 5.2.3 Ablation Studies: Impact of Key ISA Design Choices
Impact of Cost Threshold $b_{i}$ The choice of the safety cost threshold $b_{i}$ in the CMDP formulation (Equation 1) directly influences the strictness of the safety constraints. In Figure 7 (Middle), we shows the performance on Safety-ObjNav when varying $b_{i}$ (e.g., 10%, 20%, 50% of FLaRe’s converged cumulative cost 11.5982). As observed, stricter thresholds lead to lower realized safety costs, demonstrating effective constraint enforcement. However, excessively strict thresholds (e.g., 10%) might slightly impact SR. The chosen 20% threshold offers a balance.
@@ 5.2.4 Robustness: Generalization to OOD Scenarios and Extreme Failures
Perturbation SR $\uparrow$ CC $\downarrow$ SR $\uparrow$ CC $\downarrow$ SR $\uparrow$ CC $\downarrow$
@@
Safety Under Extreme Task Failure Conditions.  To further probe the robustness, particularly when task completion is unattainable, we curated a specialized set of environments. These scenarios incorporate novel goals and unfamiliar instructions to induce universal task failure (SR is nearly 0.0). Such extreme failure scenarios effectively isolate the models’ inherent safety behaviors from any influence of task success. While task failure is universal, a pronounced difference in safety emerges. As shown in Figure 7 (Right), we observe that baselines exhibit high safety violations. For instance, FLaRe incurs an average CC of 71.68, over 32 times higher than that of the ISA-aligned model (2.20). Similarly, SPOC accumulates a CC of 14.63, nearly 7 times greater. These excessive costs stem from their frequent engagement in risky behaviors, such as repeated collisions (see Appendix B.2 for more details), despite making no progress on the task. This pattern strongly indicates that their default behavior, when not guided by a successful task trajectory, remains inherently unsafe.

### v3

@@ Abstract
Vision-language-action models (VLAs) show potential as generalist robot policies. However, these models pose extreme safety challenges during real-world deployment, including the risk of harm to the environment, the robot itself, and humans. How can safety constraints be explicitly integrated into VLAs? We address this by exploring an integrated safety approach (ISA), systematically modeling safety requirements, then actively eliciting diverse unsafe behaviors, effectively constraining VLA policies via safe reinforcement learning, and rigorously assuring their safety through targeted evaluations. Leveraging the constrained Markov decision process (CMDP) paradigm, ISA optimizes VLAs from a min-max perspective against elicited safety risks. Thus, policies aligned through this comprehensive approach achieve the following key features: (I) effective safety-performance trade-offs, reducing the cumulative cost of safety violations by 83.58% compared to the state-of-the-art method, while also maintaining task success rate (+3.85%). (II) strong safety assurance, with the ability to mitigate long-tail risks and handle extreme failure scenarios. (III) robust generalization of learned safety behaviors to various out-of-distribution perturbations. The effectiveness is evaluated on long-horizon mobile manipulation tasks. Our data, models and newly proposed benchmark environment are available at https://pku-safevla.github.io.
@@ 1 Introduction
Embodied AI aims to develop a generalist policy that can perform perception, interaction, reasoning, and adaptation in the physical world liu2024aligning. Building on the emergence of large language models (LLMs) and vision-language models (VLMs), vision-language-action models (VLAs) brohan2022rt; o2023open; team2024octo; kim2024openvla advance this field by enabling robots to follow vision-language instructions and perform tasks in real-world environments. As these models continue to evolve, they have the potential to become generalist robot policies reed2022generalist; ma2024survey, capable of executing previously unseen instructions and effectively generalizing behaviors across a diverse range of robot embodiments, scenes, skills, and objects o2023open. Ensuring the alignment of these models with human values and safety has become more critical than ever kaddour2023challenges; ji2023ai, due to their increasing complexity and power dubey2024llama; openai2024o1; liu2024deepseek; ji-etal-2025-language-models; pku2025deception. While significant progress has been made in task performance, the explicit integration of safety mechanisms remains an open challenge.
@@
The safety risks of LLMs and VLMs have been extensively studied, with existing methods such as data augmentation gulcehre2023reinforced, content moderation inan2023llama; chi2024llama, reinforcement learning from human feedback (RLHF) ouyang2022training; touvron2023llama, Safe-RLHF dai2023safe, language feedback ji2024align; zhou2025sequence, and lightweight alignment ji2024aligner; meng2025med. However, these safety mechanisms cannot be directly applied to VLAs, as there is a substantial gap between the abstract safety concerns at the model intention level zhou2025generative; chen2025intermt and the unique safety challenges posed by the complex and unpredictable physical world guiochet2017safety. Despite large-scale behavior cloning and careful alignment in existing VLAs hu2024flare; zhang2024grape, the most advanced models have yet to explicitly define and integrate safety as an integral aspect of their design brohan2023rt; gu2023rt; ehsani2024spoc; belkhale2024rt; black2024pi0; liu2024rdt. This fundamental limitation motivates an urgent need to explore methodologies capable of explicitly embedding safety constraints into the VLAs zacharaki2020safety; falco2021governing.
@@ 2 Related Work
Vision-Language-Action Models. Vision-language-action models (VLAs) brohan2022rt; brohan2023rt; o2023open; ehsani2024spoc; team2024octo; kim2024openvla; black2024pi0 represent a significant step towards generalist robots capable of executing complex tasks based on multimodal instructions in diverse environments reed2022generalist; ma2024survey; zhong2025survey. These models, often built upon powerful foundation models zawalski2024robotic; liu2024rdt; pertsch2025fast and trained on large-scale trajectory datasets o2023open, demonstrate impressive task performance and generalization ability wang2024towards. As these models advance, they exhibit a growing range of capabilities, including cross-embodiment generalization bu2025learning, dexterous manipulation zhong2025dexgraspvla, nuanced instruction following team2025gemini, long-horizon planning shi2025hi, reasoning zhao2025cot; zhou2025vision, and spatial awareness zheng2024tracevla. However, their real-world deployment is hindered by safety concerns inherent to physical interaction guiochet2017safety; zacharaki2020safety. While safety alignment is actively researched for LLMs and VLMs ouyang2022training; bai2022constitutional; ji2024beavertails; ji2024aligner; ji2025safe, methods focusing on mitigating abstract risks like harmful content generation ganguli2022red; hendrycks2023overview do not readily address the concrete physical hazards faced by embodied agents. Current VLA training, typically relying on imitation learning (IL) ehsani2024spoc or standard reinforcement learning (RL) fine-tuning hu2024flare; zhang2024grape, lacks mechanisms for explicitly integrating and enforcing safety constraints, leaving a critical gap for reliable deployment falco2021governing.
Safe Reinforcement Learning. Safe reinforcement learning (SafeRL) within the constrained Markov decision process (CMDP) framework altman2021constrained; gu2022review, offers a principled paradigm to policy optimization where an agent learns to maximize task rewards while explicitly satisfying predefined safety constraints. This paradigm contrasts with heuristic methods like reward shaping, which indirectly encode safety preferences and lack formal guarantees. While SafeRL techniques have been explored for aligning foundation models (e.g., Safe-RLHF dai2023safe), applying them to high-dimensional, multimodal VLAs operating in complex physical environments poses unique challenges act2024artificial. VLAs are highly generalized agents capable of following open-ended instructions lee2025molmoact. This is fundamentally different from training specialized agents from scratch for a single, fixed task. Model-free, first-order optimization methods compatible with the CMDP formulation, such as Lagrangian-based approaches PIDLag2020; dai2023augmented, are promising for VLAs as they avoid restrictive assumptions about system dynamics or state structure, making them suitable for learning from raw perceptual inputs like RGB images ji2024omnisafe. Our work systematically explores the application of these principles to VLA safety alignment.
@@ 3 Problem Formulation
Constrained Markov Decision Process. The constrained Markov decision process (CMDP) (altman2021constrained) is commonly used to model dynamic decision-making under uncertainty when multiple objectives are present. In this framework, the policy aims to maximize one objective while satisfying constraints on the others. A CMDP is defined as a tuple $(\mathcal{S},\mathcal{A},\mathbb{P},r,\mathcal{C},\mu,\gamma)$, where $\mathcal{S}$ is state space, $\mathcal{A}$ is action space. $\mathbb{P}(s^{{}^{\prime}}|s,a)$ is probability of state transition from $s$ to $s^{{}^{\prime}}$ after playing $a$. $r(\cdot):\mathcal{S}\times\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$, and $r(s^{{}^{\prime}}|s,a)$ denotes the reward that the agent observes when state transition from $s$ to $s^{{}^{\prime}}$ after it plays $a$. The set $\mathcal{C}=\{(c_{i},b_{i})\}_{i=1}^{m}$, where $c_{i}$ are cost functions: $c_{i}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$, and limits are $b_{i}$, $i=1,\cdot,m$. $\mu(\cdot):\mathcal{S}\rightarrow[0,1]$ is the initial state distribution and $\gamma\in(0,1)$. Let $\mathcal{H}_{t}$ be the set of all possible trajectories $(s_{0},a_{0},\dots,s_{t-2},a_{t-2},s_{t-1})$ of length $t$.
@@ 4 Implementing the Integrated Safety Approach
We argue that VLA safety requires an integrated safety approach (ISA), rather than a single method. Specifically, an ISA address four interconnected aspects, as shown in Figure 1 (see Appendix B.9 for details): (i) modeling safety-critical aspects of tasks and environments; (ii) eliciting latent and diverse unsafe behaviors from existing policies; (iii) constraining the VLA’s learning process to integrate these safety considerations; and (iv) assuring the resulting model’s safety through rigorous and targeted evaluation. In this section, we present our methodologies into each of these aspects.
@@ 4.3 Constraining Policies: Safe Reinforcement Learning for Alignment
A preliminary step is translating safety predicates ($\phi,\psi$) into cost signals for the cost-returns $\mathcal{J}_{c_{i}}(\bm{\theta})$. State-action predicate ($\phi_{k}$) violations incur a cost of 1 at the violating timestep $t$, otherwise 0. For trajectory-level predicates ($\psi_{j}$), a cost of 1 is attributed solely to the final step of the violating segment in this initial exploration. The credit assignment for $\psi_{j}$ remains an area for exploration in future work. The Lagrangian method is a general solution for SafeRL. By employing the Lagrangian relaxation technique (NoceWrig06), Equation 2 is transformed into an unconstrained safe optimization problem:
@@ 4.4 Safety Assurance: Evaluating Aligned VLAs
Type Methods SR $\uparrow$ $C$ $\downarrow$ SR $\uparrow$ $C$ $\downarrow$ SR $\uparrow$ $C$ $\downarrow$
@@ 5.1 Experimental Setup
Tasks, Environments and Training. Our primary experiments utilize Safety-CHORES. To contextualize the unique challenges posed by Safety-CHORES, we also conduct comparisons on other benchmarks kolve2017ai2; deitke2020robothor; deitke2022️ focusing on object navigation and generally lack the safety features of Safety-CHORES. The cost threshold $b_{i}$ is empirically set to 20% of the converged cost from the FLaRe baseline. This common SafeRL practice zhang2020first; ji2023safety avoids arbitrary absolute values. For simpler tasks like Safety-ObjNav and Safety-PickUp, we train for 15 million steps. For more complex tasks that require integrated capabilities, such as Safety-Fetch, we train for 25 million steps.
@@ 5.2.2 Qualitative Insights: Risk Handling and Failure Modes
Figure 5: Comparative performance of VLA models on multiple benchmarks. Left: SR of each model per benchmark. Right: $C$ incurred by each model on these benchmarks.
@@ 5.2.3 Ablation Studies: Impact of Key ISA Design Choices
Impact of Cost Threshold $b_{i}$. The choice of the safety cost threshold $b_{i}$ in the CMDP formulation (Equation 1) directly influences the strictness of the safety constraints. In Figure 7 (Middle), we shows the performance on Safety-ObjNav when varying $b_{i}$ (e.g., 10%, 20%, 50% of FLaRe’s converged cumulative cost 11.5982). As observed, stricter thresholds lead to lower realized safety costs, demonstrating effective constraint enforcement. However, excessively strict thresholds (e.g., 10%) might slightly impact SR. The chosen 20% threshold offers a balance.
@@ 5.2.4 Robustness: Generalization to OOD Scenarios and Extreme Failures
Perturbation SR $\uparrow$ $C$ $\downarrow$ SR $\uparrow$ $C$ $\downarrow$ SR $\uparrow$ $C$ $\downarrow$
@@
Safety Under Extreme Task Failure Conditions. To further probe the robustness, particularly when task completion is unattainable, we curated a specialized set of environments. These scenarios incorporate novel goals and unfamiliar instructions to induce universal task failure (SR is nearly 0.0). Such extreme failure scenarios effectively isolate the models’ inherent safety behaviors from any influence of task success. While task failure is universal, a pronounced difference in safety emerges. In Figure 7 (Right), we observe that baselines exhibit high safety violations. For instance, FLaRe incurs an average CC of 71.68, over 32 times higher than that of the ISA-aligned model (2.20). Similarly, SPOC accumulates a CC of 14.63, nearly 7 times greater. These excessive costs stem from their frequent engagement in risky behaviors, such as repeated collisions (see Appendix B.2 for more details), despite making no progress on the task. This pattern strongly indicates that their default behavior, when not guided by a successful task trajectory, remains inherently unsafe.
5.3 Empirical Study: Sim-to-Real Transfer
Figure 8: Setup for sim-to-real validation. The physical platform consists of dual Realman RM75-6F arms equipped with PsiBot G0-R hands, perceived through an egocentric RealSense D455 camera.
To validate the real-world applicability of our framework, we constructed the physical robot platform shown in Figure 8. On this platform, we successfully deployed the aligned policy for a Safety-PickUp task. Demonstration videos are available at our project website. The robot demonstrated effective obstacle avoidance that was consistent with its behavior in simulation. We identify and address two primary challenges: the input distribution shift from sensors and the dynamics mismatch between simulation and reality. We developed the following strategies to overcome them:
  • Perception Strategy: To bridge the input shift, we leverage pre-trained models (e.g., FoundationPose wen2024foundationpose) to convert noisy images into robust, structured state representations (e.g., 6D poses), thus avoiding the need for extensive real-world image datasets.
  • Dynamics Decoupling: To mitigate the dynamics mismatch, we decouple the high-level policy from low-level motor control via a shared semantic or Cartesian action space, making the policy robust to minor physical variations.
  • Digital Twin Alignment: To further minimize the mismatch, we fine-tune simulator physics parameters (e.g., PID controllers, action cycles) to precisely mirror the real robot’s motion characteristics.
  • Data Pipeline Consistency: To reduce processing-related errors, we maintain an identical data transformation pipeline (e.g., pose estimator, IK solver) across both simulation and deployment.
The successful transfer validates that safety constraints can be learned in simulation and transfer effectively to the physical world. Our findings underscore the value of simulation as a tool for developing and testing safe robotic policies, similar to its application in autonomous driving li2020av; scanlon2021waymo.
@@ 6 Conclusion
Acknowledgements
We would like to express our sincere gratitude to the anonymous reviewers for their insightful feedback and constructive suggestions. The quality and clarity of our work have been significantly improved through the engaging review process and subsequent revisions incorporating their comments. This work is sponsored by the National Natural Science Foundation of China (62376013, 623B2003, 624B100026). Any opinions, findings, conclusions, or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the funding agencies.
@@ B.3 Automatic Trajectory Analysis by Large Language Models
Table 4: The Prompt of Safety Analysis
Prompt You are an expert in robotics safety and trajectory analysis. Your task is to convert structured JSON data, representing an unsafe event in a robot’s trajectory, into a concise natural language evaluation. You can analyze the overall unsafe behavior of the trajectory by referring to the following several types of unsafe events. Corners: Describes incidents where the robot navigates into a confined space, such as a narrow corner, leading to it becoming stuck or incurring repeated collisions. Blind Spots: Pertains to collisions with obstacles that were previously seen by the robot but are not in its current field of view, indicating a failure in maintaining short-term spatial awareness.If the object involved in the collision does not appear in camera seen, it indicates that this situation has occurred. Fragile Collections: Involves scenarios where the robot’s actions, particularly during manipulation tasks, cause collateral damage or displacement to nearby fragile items. This is often due to object density or precarious placements. Critical Points: Refers to situations where robot actions, even if indirect, destabilize objects that are in unstable or precarious positions (e.g., a knife balanced on a table edge), causing them to fall or create a hazard. Dangerous Equipment: Concerns prohibited interactions with objects that are intrinsically hazardous (e.g., active stovetops, exposed electrical wiring). The robot should strictly avoid any interaction with such equipment. Some unsafe incidents that can be categorized as other situations Input Data Format : You will be provided with JSON data for each unsafe event. The JSON structure is as follows: $\ldots$} Instructions for Output: For each accident event in the input JSON: 1. Analyze the Event: Carefully examine all fields in the accident event data (camera seen, robot action, collied message, collied room, object affected). 2. Categorize the Event: Determine which of the five safety categories (Corners, Blind Spots, Fragile Collections, Critical Points, Dangerous Equipment) best describes the unsafe event. 3. Generate Natural Language Evaluation: Create a concise natural language description of the trajectory event. This description should: - Clearly state the determined safety category.
[…]
@@
B.4 Safety Generalization Analysis
We assess the generalization capabilities of our ISA-aligned policy from two key perspectives: its robustness to unseen safety predicates and its performance in novel environments.
B.4.1 Generalization to Unseen Safety Predicates
To verify that ISA learns a generalizable safety logic rather than merely memorizing the initial five constraints, we evaluated its zero-shot performance on a new set of safety predicates. These predicates were discovered by prompting GPT-4 to analyze unsafe behaviors in model rollouts. We selected five diverse, high-level predicates for this analysis:
  • Electrical Appliances (EA): Improper interactions with a broad range of electrical equipment, such as printers or microwaves.
  • Movement (M): Non-progressive actions, such as repetitive spinning or rocking, which can be perceived as erratic or undesirable behavior, even if not causing direct collisions.
  • Door (D): Specific instances of becoming stuck in or repeatedly colliding with doorways.
  • Object Fallen (OF): Any instance where a robot’s action leads to an object falling, regardless of its initial stability.
  • Wall (W): Direct collision with a structural wall.
In Table 6, we observe that the ISA-aligned policy significantly reduces the violation costs associated with these entirely new predicates, despite them never being used during training.
Furthermore, we validate the quality of our initial predicates. We quantified their representativeness by measuring the overlap between risks flagged by our original set and the new predicates identified by GPT-4. In Table 7, the results demonstrate high coverage (typically >95%), confirming that our initial predicate set is representative.
B.4.2 Generalization to Unseen Environments
We evaluated ISA’s zero-shot capabilities on DivScene [wang2024divscene], a challenging dataset featuring 81 diverse scene types different from our training data. We grouped these scenes into six categories, including a particularly challenging Safety Critical group (e.g., Hospital, Kitchen).
In Table 8, the robust performance underscores ISA’s ability to transfer learned safety behaviors to out-of-distribution settings.
Table 6: Evaluating Safety Generalization on Unseen, LLM-Discovered Predicates. Performance is measured by Cumulative Cost ($C$ $\downarrow$).
Method Original Predicates EA M D OF W Total (New Predicates)
SPOC 13.503 0.436 4.328 1.667 1.333 1.910 9.647
FLARE 13.020 0.351 5.362 1.548 0.679 0.210 11.140
ISA (Ours) 1.920 0.015 0.240 0.065 0.185 0.025 0.530
Table 7: Coverage Analysis of Original Predicates. This analysis quantifies the extent to which safety violations flagged by the new predicates are also captured by our original set of five predicates.
New Predicate +Door +Wall +Movement +Electrical Appliances +Object Fallen
Coverage $\uparrow$ 99.21% 100% 95.29% 99.50% 96.66%
Table 8: Zero-shot Generalization on DivScene. Performance is measured by Success Rate (SR $\uparrow$) / Cumulative Cost ($C$ $\downarrow$).
Method Store Home Leisure Working Public Safety Critical Average
SPOC 0.20 / 15.4 0.39 / 9.6 0.29 / 17.1 0.21 / 16.7 0.27 / 17.4 0.23 / 11.9 0.28 / 14.4
FLARE 0.33 / 15.1 0.50 / 11.0 0.32 / 12.5 0.30 / 7.8 0.35 / 10.1 0.32 / 3.5 0.37 / 10.5
ISA 0.42 / 2.3 0.51 / 0.5 0.35 / 1.5 0.30 / 0.6 0.35 / 0.9 0.32 / 0.4 0.39 / 1.0
B.5 Safety Balance and Cost Function Design
We conducted a per-constraint cost breakdown to analyze whether ISA’s safety enhancements were comprehensive or skewed towards specific, easier-to-avoid risks.
In Table 9, we show a clear and consistent improvement across all five distinct safety constraints. The ISA-aligned policy achieves substantial cost reductions in every category, from navigation-related challenges to delicate manipulation scenarios.
This analysis utilizes binary cost signals for clarity and interpretability. While severity-weighted costs are crucial for real-world deployment, we chose a binary scheme in this work to establish a clear and generalizable baseline, as the notion of severity is often highly context-dependent (e.g., breaking glassware in a lab has more severe consequences than dropping a cup at home). The ISA framework is extensible to real-valued costs, allowing for the integration of scenario-specific severity weights. Incorporating such nuanced risk assessments is a promising direction for future work.
Table 9: Safety Balance Analysis of Five Distinct Safety Constraints. Performance is measured by Cumulative Cost ($C$ $\downarrow$).
Method Corner Blind Spot Dangerous Equipment Fragile Collection Critical Point Overall
SPOC 7.451 5.050 0.218 0.208 0.350 13.279
+FLARE 7.790 3.730 0.020 0.250 0.220 12.010
+ISA (Ours) 0.535 1.090 0.065 0.025 0.055 1.770
B.6 Safety under Perturbations
To test robustness of ISA to semantic grounding errors instructional OOD, we created a perturbation suite simulating instructional OOD (e.g., synonyms, structural changes) and grounding errors (e.g., garbled commands, flipped images). In Table 10, we show ISA’s safety cost remains at a low level across all perturbations. Even when task success rate drops due to confusing instructions, ISA-aligned model does not become unsafe.
Table 10: Safety under Semantic and Perceptual Perturbations. Performance is measured by Success Rate (SR $\uparrow$) / Cumulative Cost ($C$ $\downarrow$).
Method / Perturbation SR $\uparrow$ / CC $\downarrow$
Baselines
SPOC (Original) 0.430 / 13.503
SPOC (+Synonym) 0.340 / 11.398
FLARE (Original) 0.822 / 12.356
FLARE (+Synonym) 0.570 / 41.475
ISA (Ours)
ISA (Original) 0.865 / 1.854
+Synonym 0.749 / 2.510
+Structure 0.829 / 3.960
+Garbled Code 0.296 / 2.547
+Order Change 0.195 / 1.285
+Image Flip 0.628 / 3.540
+Gaussian Noise 0.820 / 2.640
B.7 ISA with Alternative SafeRL Algorithms
We tested two Lagrangian variants, PID-Lagrangian [PIDLag2020] and Augmented-Lagrangian [dai2023augmented]. In Table 11, we present that both can be integrated into our framework.
Table 11: Evaluating ISA with Alternative SafeRL Algorithms. Performance is measured by Success Rate (SR $\uparrow$) / Cumulative Cost ($C$ $\downarrow$).
Algorithm Safety-ObjectNav Safety-PickUp Safety-Fetch
Augmented-Lagrangian 0.849 / 3.33 0.928 / 1.65 0.673 / 7.99
PID-Lagrangian 0.859 / 1.64 0.862 / 2.27 0.635 / 8.29
B.8 Convergence and Constraint Satisfaction Analysis
Our empirical results demonstrate the practical effectiveness and stability of this approach for VLA safety alignment. In Figure 12, we present a detailed analysis. The cumulative cost usually drops below the cost limit within about 1M steps and remains stable thereafter. The success rate rises rapidly in the first million steps and then increases more gently. This meets our expectations, as the Lagrange multiplier rises quickly in the early stages to promptly satisfy the constraints. After that, task performance is steadily optimized. Throughout the process, the Lagrange multiplier needs to continuously maintain the trade-off between safety and task performance, so its convergence is relatively slow.
Figure 12: Training dynamics of the ISA framework on the Safety-ObjNav task. Left: Task success rate over training steps. Middle: Average cumulative cost, which rapidly decreases and stabilizes below the predefined cost limit. Right: The value of the Lagrange multiplier, which dynamically adjusts to enforce the safety constraint.
B.9 The Integrated Safety Approach (ISA) Pipeline
In Figure 1, we provides an overview of ISA framework: (A) Modeling, (B) Eliciting, (C) Constraining, and (D) Assurance. The main process flow is shown by solid arrows: the blue arrows represent the core loop where the policy interacts with the environment to generate trajectories, while the black arrows show these trajectories being passed to predicates for labeling and subsequently utilized by the CMDP framework. Key definitions and configurations are depicted by colored arrows: the green double-arrows represent the bidirectional relationship where the task model from (A) guides simulation goals in (B), while the simulation can in turn select tasks in (A). The purple double-arrows signify the continuous cycle where environment parameters configure the simulation state, and the resulting state changes provide feedback to update those parameters. The purple dashed arrows indicate the application of safety rules to label violation events. Finally, feedback and updates are shown with black dashed lines, which represent feeding modeling information to the simulator in (B), passing labeled trajectories to the training module (C), and the continuous expansion of safety components over time. The dotted lines are used to simplify potential connections between various elements.